Circular No. 7157

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 H4 (SOHO) D. A. Biesecker, SM&A Corporation and Goddard Space Flight Center,

reports observations of another apparent Kreutz sungrazing comet that was discovered by S. Gregory (University of Maryland) and was detected by both the C2 and C3 coronagraphs. There was a short tail evident, and on Apr. 27.6 UT the comet’s magnitude was found to be ∼ 4.6. Astrometric measurements, made by Gregory and Biesecker (and reduced by G. V. Williams, Smithsonian Astrophysical Observatory), together with orbital elements by B. G. Marsden, are given on MPEC 1999-J05.

1999 UT  α_{2000}  δ_{2000}
Apr. 26.821  2^{h}29.^{m}7  +11^{o}29′

SUPERNOVA 1999bz IN UGC 8959 M. Modjaz and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report their discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT), of an apparent supernova in UGC 8959. SN 1999bz was discovered on an unfiltered image taken on May 1.4 UT (mag ∼ 17.6). Inspection of an earlier KAIT image on Apr. 24.4 shows a hint of the new star. The supernova is located at α = 14^{h}01^{m}46^{s}.75, δ = +68^{o}52′46′′.5 (equinox 2000.0), which is ∼ 12′′ west and 8′′ north of the nucleus of UGC 8959. A KAIT image of the same field on Apr. 19.4 (limiting mag ∼ 19) showed nothing at the position of SN 1999bz.

SUPERNOVA 1999by IN NGC 2841 CCD magnitudes, unfiltered unless otherwise noted: Apr. 30.826 UT,

15.5 (Ernő Berkó, Ludányhalászi, Hungary; independent discovery, 0.1-m refractor); May 1.885, 14.5 (M. Armstrong, Rolvenden, Kent, England); 2.872, V = 14.42 ± 0.05, B − V = +0.27 ± 0.05, V − I = +0.11 ± 0.02 (D. Hanžl, Brno, Czech Republic; comparison star GSC 3431.625, using Tycho magnitudes V = 11.05, B − V = +0.564, V − I = +0.64).

V4444 SAGITTARII Visual magnitude estimates: Apr. 30.644 UT, 9.1 (A. Pearce, Ned-

lands, W. Australia); May 1.640, 9.6 (Pearce); 2.170, 10.0 (T. A. Napoleao, São Paulo, Brazil); 3.052, 9.5 (A. Baransky, Kiev, Ukraine).

1999 May 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green